OpenStudy (anonymous):

A ball is dropped from rest from a height of 200m. What is the height and velocity of the ball 3 seconds later? Is this calc-based physics? If so, please walk me through it.

OpenStudy (anonymous):

its not...

OpenStudy (anonymous):

It's in my friends Calc 1 book, that's the only reason why I ask.

OpenStudy (anonymous):

its easy just use these equatioons: v=u+at; s=ut+1/2at^2 v^2=u^2+2as

OpenStudy (anonymous):

And u stands for....

OpenStudy (anonymous):

initial velocity=u

OpenStudy (anonymous):

@KenField

OpenStudy (anonymous):

from rest means u=0

OpenStudy (anonymous):

You have an expression for acceleration (g), integrate for velocity after a dt of 3 seconds, now you have v, integrate again for dx

OpenStudy (anonymous):

to solve this in calc (though unnecessary, but ah well, the heck with it), we'll start with the following. acceleration is defined as: \(\vec a = \frac{d \vec v }{d t}\) rearraging, \( d \vec v = \vec a dt\) integrating both sides with respect to t, \(\vec v =\vec a t +C\) since when \(t=0\), \(\vec v=0\), C=0. but when \(t=3\), \(\vec a =-9.8 a_y\) where \(a_y\) is the unit vector in the y direction. so, \(\vec v =-29.4a_y\) now remember that \(\vec v =\frac{d \vec r}{dt}\) integrating in the same way with acceleration, \(\vec r = \vec v t +C\), where\( C=200a_y\) when t=0 and r=0. now, substituting the values for \(\vec v_{t=3}\) and t=3, \(\vec r = -9.8 a_y(3) +200a_y=170.6 a_y\) is there a part where you need further help?

OpenStudy (anonymous):

correction: \(\vec r=\vec v t+C\), where \(C=200a_y\) when t=0

OpenStudy (shane_b):

Here's the easiest way in my opinion: \[\large \text{distance fallen in 3 seconds}= \frac{1}{2}gt^2=\frac{1}{2}(-9.8m/s^2)(3.0s)=-44.1m\]\[\large \text{height after 3 seconds}=200m-44.1m=155.9m\]\[\large \text{velocity after 3 seconds}=at=(-9.8m/s^2)(3.0s)=-29.4m/s\]